A corporate bond's face value is $1000 and it matures in 8 years. Coupons are paid semi-annually. What is the bond's yield to maturity, if its coupon rate is 6.4% and it is currently trading for $1080.9? [Provide your answer in percent, with two decimals, omitting the % sign.]
Answer #1

Number of periods = 8 * 2 = 16

Coupon = (0.064 * 1000) / 2 = 32

Yield to maturity = 5.15

Keys to use in a financial calculator:

2nd I/Y 2

FV 1000

PV -1080.9

PMT 32

N 16

CPT I/Y